# Phase Transition of Spectral Fluctuations in Large Gram Matrices with a Variance Profile: A Unified Framework for Sparse CLTs

### v1

@@ 3.3 CLT in the high-sparsity regime
We now turn to the CLT for the LSS in the high-sparsity regime, where $s\to 0$ and $q^{2}=n^{O(1)}$. Unlike the dense or moderate-sparsity settings, the asymptotic mean term in Theorem 3.3 diverges as $s\to 0$. This divergence arises from a mismatch in convergence rates: the variance of the LSS scales as $\sqrt{p}\,q$, while the mean scales as $q^{2}$. Consequently, when $q^{2}\ll\sqrt{p}\,q$, the standard centralization fails to remove the dominant deterministic component, leading to a degenerate limit.
@@
$L^{c}\left(f,\mathbf{S}_{n}\right)=\sqrt{p}q\int f(x)\mathrm{d}\left[F^{\mathbf{S}_{n}}(x)-\pi_{n}(x)\right]+\frac{\sqrt{p}q}{2\pi\mathbf{i}}\oint_{\mathcal{C}_{1}}f(z)\mathscr{E}_{n}(z)\mathrm{d}z$
where $\mathscr{E}_{n}(z)=n^{-1}\sum_{j=1}^{n}\psi_{j,n}(z)$, and $\psi_{j,n}(z)$ are the unique solutions to the system of linear equations
$\psi_{j}(z)=\sum_{m=1}^{n}a_{jm}\,\psi_{m}(z)+\theta_{j}(z),\qquad 1\leq j\leq n.$
@@
Let $f_{1},\dots,f_{k}$ be functions that are analytic in a complex neighborhood of the support of the LSD of $\mathbf{S}_{n}$, and continuous at zero. Under Assumptions 2.12.4 with $1/4<\phi<1/2$, the random vector $\left(L^{c}(f_{j},\mathbf{S}_{n})\right)_{j=1}^{k}$ is tight and converges weakly to a Gaussian vector $\left(X_{f_{j}}\right)_{j=1}^{k}$, where the means $\mu_{n}(X_{f})=0$ and covariance functions are defined as follows:
@@
To further illustrate this point, we conduct a simulation at the boundary case $\phi=1/4$. The model follows the sparse MIMO channel formulated in (6) of Section 4.2, where the diagonal entries are generated as $d_{i}\sim\mathrm{Uniform}(6,8)$ for $i=1,\ldots,N_{r}$ and $\tilde{d}_{j}\sim\mathrm{Uniform}(6,8)$ for $j=1,\ldots,N_{t}$. The matrix $\mathbf{X}$ has i.i.d. entries $x_{ij}\sim\mathcal{CN}(0,1)$, and $\sigma^{2}$ is set to 1. Throughout, we take $(N_{r},N_{t})=(5000,10000)$ and average over $2,000$ independent channel realizations. The statistic considered is $\widetilde{T}_{\log}={q}/{\sqrt{N_{r}}}\bigl\{C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})-V(\sigma^{2})\bigr\}-\mu_{\log}$, where $C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})$ and $\mu_{\log}$ are defined in Theorem 4.5. It is centered so that $\widetilde{T}_{\log}$ converges to $\mathcal{N}(0,\sigma_{\log}^{2})$. For fixed $n$, reducing $q$ enhances the magnitude of the mean deviation. This trend is clearly visible when comparing the cases $q=0.7n^{1/4}$, $q=0.6n^{1/4}$, and $q=0.5n^{1/4}$, where the smaller values of $q$ lead to more pronounced mean drift (Figure 1). This behavior indicates that the $O(\sqrt{n}/q^{2})$ term remains non-negligible when $q\asymp n^{1/4}$. Its contribution is sufficiently large to distort the mean, which in turn motivates the restriction $1/4<\phi<1/2$, under which a clean asymptotic expression can be established.
@@
The combination of Theorem 3.3 and Corollary 3.5 highlights two distinct phenomena concerning the asymptotic fluctuations of the LSS. (1) Phase transition. As the sparsity parameter $s$ decreases to zero, the limiting law in Corollary 3.5 emerges as a continuous extension of Theorem 3.3, exhibiting a clear phase transition between moderate and high sparsity. While the LSD remains unchanged, the fluctuation structure shifts from the moderate-sparsity regime to the high-sparsity regime, illustrating how sparsity affects the second-order limiting behavior. (2) Fourth-moment-driven fluctuation. In the high-sparsity regime, the asymptotic mean and variance are governed by the term originating from the fourth-moment structure of the entries, which appears with coefficient $\tilde{\nu}_{4}$ after taking the sparsity limit. While the resulting expressions still involve the diagonal second-moment structure through $\widetilde{\mathbf{\Sigma}}_{i}^{2}$ and $\mathbf{\Sigma}_{j}^{2}$, the components that are originally scaled by $s$ are suppressed. By contrast, under moderate sparsity, multiple fluctuation contributions coexist and jointly determine the asymptotic mean and variance. This highlights a distinct fluctuation regime induced by high sparsity.
Taken together, Theorem 3.3 and Corollary 3.5 establish a unified CLT framework for LSS of large sparse Gram matrices. This framework encompasses a broad range of sparsity regimes, from moderate sparsity to high sparsity, and applies to both Gaussian and non-Gaussian distributions, as well as to real and complex-valued entries. It thereby provides a comprehensive and robust foundation for analyzing second-order fluctuations in high-dimensional models.
@@ 4.1 Application to equality test of two large-scale fading matrices
where $\mathbb{B}=(b_{ij})$ has independent entries $b_{ij}\sim{\rm Bernoulli}(1,s)$ with sparsity level $s=q^{2}/N_{t}$. The matrix $\mathbf{L}=(l_{ij})$ contains the large-scale fading coefficients, whose the slow variation reflects the macroscopic propagation environment. The matrix $\mathbf{G}=(g_{ij})$ models i.i.d. small-scale fading, typically circular Gaussian $g_{ij}\sim\mathcal{CN}(0,1)$, though more general distributions with finite moments can also be accommodated.
@@
$\mathbf{H}_{1}\in\mathbb{R}^{p\times n},\quad\mathbf{H}_{2}\in\mathbb{R}^{p\times n},$
@@
Define the sample covariance matrices as
@@
the trace difference $\operatorname{Tr}(\mathbf{S}_{1})-\operatorname{Tr}(\mathbf{S}_{2})$ provides an unbiased estimate of the large-scale energy gap $1/N\left[\operatorname{Tr}(\mathbf{L}_{1}\mathbf{L}_{1}^{*})-\operatorname{Tr}(\mathbf{L}_{2}\mathbf{L}_{2}^{*})\right]$. Motivated by this observation, a natural choice is the linear test function $f(x)=x$, for which the corresponding LSS reduces to the trace. Accordingly, we define the test statistic
@@
The term $1/(N_{r}N_{t})\sum_{i,j}\mathbb{E}(l_{ij,1}g_{ij,1})^{4}$ can be estimated by $sN_{t}/N_{r}\sum_{i,j}h_{ij,1}^{4}$, which is unbiased and consistent (using $\hat{s}=1/(N_{r}N_{t})\sum_{i,j}I(h_{ij,1}\neq 0)$ if $s$ is unknown). Assuming that $\mathbb{E}(g_{ij,1}^{4})=\tilde{\nu}_{4}$, the second term, ${s}/({N_{r}N_{t}})\sum_{i,j}l_{ij,1}^{4}$, can be estimated analogously by $sN_{t}/(N_{r}\tilde{\nu}_{4})\sum_{i,j}h_{ij,1}^{4}$.
Moreover, the technical restriction $\phi>1/4$ in Theorem 3.3, imposed to control an additional term of order $O(\sqrt{N_{r}}/q^{2})$ in the mean expansion, can be removed under $\mathcal{H}_{0}$. In this case, since the mean is identically zero, the potentially divergent term no longer arises, and the CLT in Theorem 4.1 remains valid for all $N_{t}^{\phi}\leq q\leq N_{t}^{1/2}$ with $\phi>0$.
@@
$\left\{D_{x}>\mu_{H_{0}}+\sqrt{2}\sigma_{H_{0}}z_{1-\alpha}\right\}$
@@
$\sigma_{H_{1}}^{2}=\frac{1}{N_{r}N_{t}}\sum_{i,j}\left[\mathbb{E}(l_{ij,1}g_{ij,1})^{4}+\mathbb{E}(l_{ij,1}g_{ij,1})^{4}\right]-\frac{s}{N_{r}N_{t}}\sum_{i,j}(l_{ij,1}^{4}+l_{ij,2}^{4}).$
@@
Two observations apply under the alternative hypothesis. First, the term $1/(N_{r}N_{t})\sum_{i,j}l_{ij,k}^{2},k=1,2$ appearing in Theorem 4.3 can be consistently estimated from $1/N_{r}\sum_{i,j}h_{ij,k}^{2},k=1,2$. Second, under $\mathcal{H}_{1}$, the restriction on $\phi$ can likewise be relaxed. For $f(x)=x$, the CLT remains valid for all $N_{t}^{\phi}\leq q\leq N_{t}^{1/2}$ with $\phi>0$.
@@ 4.2 Application to outage probability analysis in sparse MIMO channels
which is a random variable due to the randomness of the channel matrix $\mathbf{H}$. This randomness motivates us to investigate fluctuation of $C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})$, which can be used to compute outage probability.
@@
Let $z=-\sigma^{2}$. The CLT for the mutual information, $C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})$, can be expressed as
$T_{\log}=\frac{q/\sqrt{N_{r}}\left\{C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})-V(\sigma^{2})\right\}-\mu_{\log}}{\sigma_{\log}}\xrightarrow{\mathcal{D}}\mathcal{N}(0,1),$
@@
$V(\sigma^{2})=-\log\det(\sigma^{2}\mathbf{T}(z))+\log\det\!\left(\mathbf{I}+\delta\widetilde{\mathbf{D}}\right)-N_{t}\sigma^{2}\delta(z)\tilde{\delta}(z),$
$\mu_{\log}$ $=-\frac{[\tilde{\nu}_{4}-(\kappa+2)s]\sigma^{4}}{2q\sqrt{N_{r}}N_{t}}\operatorname{Tr}\!\left(\mathbf{D}^{2}\mathbf{T}^{2}(z)\right)\operatorname{Tr}\!\left(\widetilde{\mathbf{D}}^{2}\widetilde{\mathbf{T}}^{2}(z)\right),$
@@
$\sigma_{\log}=$ $\frac{\left[\tilde{\nu}_{4}-(\kappa+2)s\right]\sigma^{4}}{N_{r}N_{t}}\operatorname{Tr}\!\left(\mathbf{D}^{2}\mathbf{T}^{2}(z)\right)\operatorname{Tr}\!\left(\widetilde{\mathbf{D}}^{2}\widetilde{\mathbf{T}}^{2}(z)\right)$ (8)
@@
$P_{\rm out}(R,\rho)\ \approx\ \Phi\!\left(\frac{\displaystyle\frac{q}{\sqrt{N_{r}}}\left(R-V(\sigma^{2})\right)-\mu_{\log}}{\sigma_{\log}}\right),$
where $\Phi$ is the standard normal CDF.
@@ 5.1 Equality test of two large-scale fading matrices
$l_{ij,1}\sim\mathrm{Uniform}(2,4),\qquad l_{ij,2}=l_{ij,1}-\theta,$
where $\theta$ controls the deviation between the two environments: $\theta=0$ corresponds to the null hypothesis $\mathcal{H}_{0}$, whereas $\theta\neq 0$ represents the alternative $\mathcal{H}_{1}$. To introduce sparsity, the missing indicators are generated as $b_{ij,1},b_{ij,2}\sim\rm Bernoulli(1,s)$, with $s=q^{2}/N_{t}$. Two regimes of $q$ are examined, corresponding to $q=N_{t}^{1/3}$ (high-sparsity regime) and $q=0.5\,N_{t}^{1/2}$ (moderate-sparsity regime). We examine two distributions for the small-scale fading terms:
@@
Tables 1 and 2 report the empirical sizes and powers of the proposed test $T_{x}$, bassed on 2,000 Monte Carlo replications, under different population distributions and system dimensions. Specifically, Table 1 corresponds to the setting $q=0.5N_{t}^{1/2}$ while Table 2 considers $q=N_{t}^{1/3}$. The results are presented for both Gaussian and Gamma populations with various ratios $c_{N}=N_{r}/N_{t}\in\{0.8,0.5,0.25\}$ and receive antenna numbers $N_{r}\in\{200,400,800\}$. The nominal significance level is fixed at $0.05$.

### v2

@@ 3.3 CLT in the high-sparsity regime
We now turn to the CLT for the LSS in the high-sparsity regime, where $s\to 0$ with $n^{\phi}\leq q<n^{1/2}$ for some fixed $\phi>0$. Unlike the dense or moderate-sparsity settings, the asymptotic mean term in Theorem 3.3 diverges as $s\to 0$. This divergence stems from a mismatch between the fluctuation and centering terms under the CLT normalization: after scaling by $\sqrt{p}q$, the variance of the LSS in (4) remains of order one, whereas the corresponding mean contribution grows at rate $\sqrt{p}/q$. Consequently, when $q^{2}\ll\sqrt{p}\,q$, the standard centering fails to remove the dominant deterministic component, leading to a degenerate limit.
@@
$L^{\tilde{c}}\left(f,\mathbf{S}_{n}\right)=\sqrt{p}q\int f(x)\mathrm{d}\left[F^{\mathbf{S}_{n}}(x)-\pi_{n}(x)\right]+\frac{1}{2\pi\mathbf{i}}\oint_{\mathcal{C}_{1}}f(z)\mathscr{E}_{n}(z)\mathrm{d}z,$
where $\mathscr{E}_{n}(z)=n^{-1}\sum_{j=1}^{n}\psi_{j}(z)$, and $\{\psi_{j}(z)\}_{j=1}^{n}$ is the unique solution to the system of linear equations
$\psi_{j}(z)=\sum_{m=1}^{n}a_{jm}(z,z)\psi_{m}(z)+\theta_{j}(z),\qquad 1\leq j\leq n.$
@@
Let $f_{1},\dots,f_{k}$ be functions that are analytic in a complex neighborhood of the support of the LSD of $\mathbf{S}_{n}$, and continuous at zero. Under Assumptions 2.12.4 with $1/4<\phi<1/2$, the random vector $\left(L^{\tilde{c}}(f_{j},\mathbf{S}_{n})\right)_{j=1}^{k}$ is tight and converges weakly to a Gaussian vector $\left(X_{f_{j}}\right)_{j=1}^{k}$, where the mean $\mu_{n}(X_{f})=0$ and the covariance function between $X_{f}$ and $X_{g}$ is defined as follows:
@@
The contours $\mathcal{C}_{1},\mathcal{C}_{2}$ are closed, non-overlapping, positively oriented, and each encloses the support of $\pi_{n}$.
@@
To further illustrate this point, we conduct a simulation at the boundary case $q\asymp n^{\phi}$ with $\phi=1/4$. The model follows the sparse MIMO channel formulated in (7) of Section 4.2, with the dimensions $(N_{r},N_{t})$ therein replaced by $(p,n)$. The diagonal entries of deterministic matrices $\mathbf{D}$ and $\widetilde{\mathbf{D}}$ are generated as $d_{i}\sim\text{Uniform}(6,8)$ for $i=1,\ldots,p$ and $\tilde{d}_{j}\sim\text{Uniform}(6,8)$ for $j=1,\ldots,n$. The matrix $\mathbf{X}$ has i.i.d. entries $x_{ij}\sim\mathcal{CN}(0,1)$, and $\sigma^{2}$ is set to 1. Throughout, we take $(p,n)=(5000,10000)$ and average over $2,000$ independent channel realizations. The statistic considered is $\widetilde{T}_{\log}={q}/{\sqrt{n}}\bigl\{C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})-V(\sigma^{2})\bigr\}-\mu_{\log}$, where $C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})$ and $\mu_{\log}$ are defined in Theorem 4.5. It is centered so that $\widetilde{T}_{\log}$ converges to $\mathcal{N}(0,\sigma_{\log}^{2})$. For fixed $n$, reducing $q$ enhances the magnitude of the mean deviation. This trend is clearly visible when comparing the cases $q=0.7n^{1/4}$, $q=0.6n^{1/4}$, and $q=0.5n^{1/4}$, where the smaller values of $q$ lead to more pronounced mean drift (Figure 1). This behavior indicates that the $O(\sqrt{n}/q^{2})$ term remains non-negligible when $q\asymp n^{1/4}$. Its contribution is sufficiently large to distort the mean, which in turn motivates the restriction $1/4<\phi<1/2$, under which a clean asymptotic expression can be established.
@@
The combination of Theorem 3.3 and Corollary 3.5 highlights two distinct phenomena concerning the asymptotic fluctuations of the LSS. (1) Phase transition. As the sparsity parameter $s$ decreases to zero, the limiting law in Corollary 3.5 emerges as a continuous extension of Theorem 3.3, exhibiting a clear phase transition between moderate- and high-sparsity. While the LSD remains unchanged, the fluctuation structure shifts from the moderate-sparsity regime to the high-sparsity regime, illustrating how sparsity affects the second-order limiting behavior. (2) Fourth-moment-driven fluctuation. In the high-sparsity regime, the asymptotic mean and variance are governed by the term originating from the fourth-moment structure of the entries, which appears with coefficient $\tilde{\nu}_{4}$ after taking the sparsity limit. While the resulting expressions still involve the diagonal second-moment structure through $\widetilde{\mathbf{\Sigma}}_{i}^{2}$ and $\mathbf{\Sigma}_{j}^{2}$, the components that are originally scaled by $s$ are suppressed. By contrast, under moderate sparsity, multiple fluctuation contributions coexist and jointly determine the asymptotic mean and variance. This highlights a distinct fluctuation regime induced by high sparsity.
Taken together, Theorem 3.3 and Corollary 3.5 establish a unified CLT framework for LSS of large sparse Gram matrices. This framework encompasses a broad range of sparsity regimes, from moderate-sparsity to high-sparsity, and applies to both Gaussian and non-Gaussian distributions, as well as to real and complex-valued entries. It thereby provides a comprehensive and robust foundation for analyzing second-order fluctuations in high-dimensional models.
@@ 4.1 Application to equality test of two large-scale fading matrices
where $\mathbb{B}=(b_{ij})$ has independent entries $b_{ij}\sim{\rm Bernoulli}(1,s)$ with sparsity level $s=q^{2}/N_{t}$. The matrix $\mathbf{L}=(l_{ij})$ contains the large-scale fading coefficients, whose slow variation reflects the macroscopic propagation environment. The matrix $\mathbf{G}=(g_{ij})$ models i.i.d. small-scale fading, typically circular Gaussian $g_{ij}\sim\mathcal{CN}(0,1)$, though more general distributions with finite moments can also be accommodated.
@@
$\mathbf{H}_{1}\in\mathbb{C}^{N_{r}\times N_{t}},\quad\mathbf{H}_{2}\in\mathbb{C}^{N_{r}\times N_{t}},$
@@
Define the Gram matrices as
@@
the trace difference $\operatorname{Tr}(\mathbf{S}_{1})-\operatorname{Tr}(\mathbf{S}_{2})$ provides an unbiased estimate of the large-scale energy gap $1/N_{t}\left[\operatorname{Tr}(\mathbf{L}_{1}\mathbf{L}_{1}^{*})-\operatorname{Tr}(\mathbf{L}_{2}\mathbf{L}_{2}^{*})\right]$. Motivated by this observation, a natural choice is the linear test function $f(x)=x$, for which the corresponding LSS reduces to the trace. Accordingly, we define the test statistic
@@
The quantity $1/(N_{r}N_{t})\sum_{i,j}\mathbb{E}(l_{ij,1}g_{ij,1})^{4}$ can be consistently estimated by $sN_{t}/N_{r}\sum_{i,j}h_{ij,1}^{4}$, which is unbiased (using $\hat{s}=1/(N_{r}N_{t})\sum_{i,j}I(h_{ij,1}\neq 0)$ if $s$ is unknown). Assuming that $\mathbb{E}(g_{ij,1}^{4})=\tilde{\nu}_{4}$, the second term, ${s}/({N_{r}N_{t}})\sum_{i,j}l_{ij,1}^{4}$, can be estimated analogously by $sN_{t}/(N_{r}\tilde{\nu}_{4})\sum_{i,j}h_{ij,1}^{4}$.
Moreover, the technical restriction $\phi>1/4$ in Theorem 3.3, imposed to control an additional term of order $O(\sqrt{N_{t}}/q^{2})$ in the mean expansion, can be removed under $\mathcal{H}_{0}$. In this case, since the mean is identically zero, the potentially divergent term no longer arises, and the CLT in Theorem 4.1 remains valid for all $N_{t}^{\phi}\leq q\leq N_{t}^{1/2}$ with $\phi>0$.
@@
$\left\{D_{x}>\mu_{H_{0}}+\sqrt{2}\sigma_{H_{0}}z_{1-\alpha}\right\},$
@@
$\sigma_{H_{1}}^{2}=\frac{1}{N_{r}N_{t}}\sum_{i,j}\left[\mathbb{E}(l_{ij,1}g_{ij,1})^{4}+\mathbb{E}(l_{ij,2}g_{ij,2})^{4}\right]-\frac{s}{N_{r}N_{t}}\sum_{i,j}(l_{ij,1}^{4}+l_{ij,2}^{4}).$
@@
Two observations apply under the alternative hypothesis. First, the terms $1/(N_{r}N_{t})\sum_{i,j}\mathbb{E}(l_{ij,k}g_{ij,k})^{4}$ and $s/(N_{r}N_{t})\sum_{i,j}l_{ij,k}^{2},k=1,2$ appearing in Theorem 4.3 can be consistently estimated in the same manner as in Theorem 4.1. Second, under $\mathcal{H}_{1}$, the restriction on $\phi$ can be relaxed as well, since for the test function $f(x)=x$, the CLT remains valid for all $N_{t}^{\phi}\leq q\leq N_{t}^{1/2}$ with $\phi>0$.
@@ 4.2 Application to outage probability analysis in sparse MIMO channels
which is a random variable due to the randomness of the channel matrix $\mathbf{H}$. This randomness motivates us to investigate the fluctuation of $C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})$, which can be used to compute outage probability.
@@
Let $z=-\sigma^{2}$ and $1/4<\phi\leq 1/2$. The CLT for the mutual information, $C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})$, can be expressed as
$T_{\log}=\frac{q/\sqrt{N_{r}}\left[C_{\mathbf{H}\mathbf{H}^{*}}(\sigma^{2})-V(\sigma^{2})\right]-\mu_{\log}}{\sigma_{\log}}\xrightarrow{\mathcal{D}}\mathcal{N}(0,1),$
@@
$V(\sigma^{2})=-\log\det(\sigma^{2}\mathbf{T}(z))+\log\det\!\left(\mathbf{I}+\delta(z)\widetilde{\mathbf{D}}\right)-N_{t}\sigma^{2}\delta(z)\tilde{\delta}(z),$
$\mu_{\log}$ $=-\frac{\left(\tilde{\nu}_{4}-(\kappa+2)s\right)\sigma^{4}}{2q\sqrt{N_{r}}N_{t}}\operatorname{Tr}\!\left(\mathbf{D}^{2}\mathbf{T}^{2}(z)\right)\operatorname{Tr}\!\left(\widetilde{\mathbf{D}}^{2}\widetilde{\mathbf{T}}^{2}(z)\right),$
@@
$\sigma_{\log}=$ $\frac{\left(\tilde{\nu}_{4}-(\kappa+2)s\right)\sigma^{4}}{N_{r}N_{t}}\operatorname{Tr}\!\left(\mathbf{D}^{2}\mathbf{T}^{2}(z)\right)\operatorname{Tr}\!\left(\widetilde{\mathbf{D}}^{2}\widetilde{\mathbf{T}}^{2}(z)\right)$ (9)
@@
$P_{\rm out}(R,\rho)\ \approx\ \Phi\!\left(\frac{\displaystyle{q}/{\sqrt{N_{r}}}\left(R-V(\sigma^{2})\right)-\mu_{\log}}{\sigma_{\log}}\right),$
where $\Phi$ is the standard normal Cumulative Distribution Function.
@@ 5.1 Equality test of two large-scale fading matrices
$l_{ij,1}\sim\text{Uniform}(2,4),\qquad l_{ij,2}=l_{ij,1}-\theta,$
where $\theta$ controls the deviation between the two environments: $\theta=0$ corresponds to the null hypothesis $\mathcal{H}_{0}$, whereas $\theta\neq 0$ represents the alternative $\mathcal{H}_{1}$. To introduce sparsity, the missing indicators are generated as $b_{ij,1},b_{ij,2}\sim\rm Bernoulli(1,s)$, with $s=q^{2}/N_{t}$. Two regimes of $q$ are examined, corresponding to $q=0.5\,N_{t}^{1/2}$ (moderate-sparsity regime) and $q=N_{t}^{1/3}$ (high-sparsity regime). We examine two distributions for the small-scale fading terms:
@@
Tables 1 and 2 report the empirical sizes and powers of the proposed test $T_{x}$ under different population distributions and system dimensions. Specifically, Table 1 corresponds to the setting $q=0.5N_{t}^{1/2}$ while Table 2 considers $q=N_{t}^{1/3}$. The results are presented for both Gaussian and Gamma populations with various ratios $c_{N}=N_{r}/N_{t}\in\{0.8,0.5,0.25\}$ and receive antenna numbers $N_{r}\in\{200,400,800\}$. The nominal significance level is fixed at $0.05$.